OpenStudy (anonymous):

Nick's restaurant bill totaled 64.20. Show how to estimate 15% of the tip. Show and explain each step

OpenStudy (anonymous):

First, x% means \[\frac{ x }{ 100 }\]

OpenStudy (anonymous):

and then 15% is like this \[\frac{ 15 }{ 100 }\]

OpenStudy (anonymous):

so you can put 64.20 in front of it and that's \[64.20*\frac{ 15 }{ 100 }\]